Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability problem

Expert replies
by auppala » Tue Mar 05, 2013 10:27 am
The probability that a teacher will give an unannounced test during any class meeting is 1/5.If a student is absent twice,then the probability that the student will miss at least one test is
A.4/5
B.2/5
C.7/75
D.9/25
E.8/25

Not sure what the official answer is. Confused between two of the options. Help please !
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Tue Mar 05, 2013 10:36 am
auppala wrote:The probability that a teacher will give an unannounced test during any class meeting is 1/5.If a student is absent twice,then the probability that the student will miss at least one test is
Probability of missing at least one test = Probability of missing test in only one day + Probability of missing test on both the day

Probability of missing test in only one day = (Probability of happening test on one of the day he was absent)*(Probability of not happening test on the other day he was absent) = (1/5)*(1 - 1/5) = (1/5)*(4/5) = 4/25

Now the above can happen in two ways : {test on first day, no test on second day} or {no test on first day, test on second day}

Hence, probability of missing test in only one day = 2*(4/25) = 8/25

And, probability of missing test on both the day = (1/5)*(1/5) = 1/25

Hence, required probability = 8/25 + 1/25 = 9/25

The correct answer is D.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by Brent@GMATPrepNow » Tue Mar 05, 2013 10:37 am
auppala wrote:The probability that a teacher will give an unannounced test during any class meeting is 1/5.If a student is absent twice,then the probability that the student will miss at least one test is
A.4/5
B.2/5
C.7/75
D.9/25
E.8/25
Questions with "at least" are often great candidates for using the complement.
That is: P(Event A occurs) = 1 - P(Event A does not occur)

So, P(miss at least 1 test) = 1 - P(miss no tests)

P(miss no tests)
P(miss no tests) = P(no test on day 1 AND no test on day 2)
= P(no test on day 1) X P(no test on day 2)
= (4/5) X (4/5)
= 16/25


So, P(miss at least 1 test) = 1 - P(miss no tests)
= 1 - 16/25
= [spoiler]9/25 = D[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by auppala » Tue Mar 05, 2013 11:17 am
Thanks! Really appreciate the quick turn around.


I used an approach similar to that used by Brent.
Join the discussion